OpenStudy (anonymous):

- 3/10 + 2/5 fractions

OpenStudy (anonymous):

-3/10+2/5 = -3/10 +4/10 = 1/10 You have to covert the second fraction so both denominators match. Least common denominator is 10. For example 2/4 = 1/2.